Terell is thinking of 2 numbers. The sum of the numbers is15. One of the numbers is 6 less than twice the other number. Which of the numbers below is the greater number?
A.7. B.8. C.14. D.16. E.30

Answer by jim_thompson5910(35256)

x+y=15
y = 2x-6


x+y=15

x+2x-6=15

3x-6=15

3x=15+6

3x=21

x = 21/3

x = 7

---------------------

y = 2x-6

y = 2(7)-6

y = 14-6

y = 8


So the smaller number is 7